40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Feb 10 17:41:25 2020, anchal, DailyProgress, scatter, Cleaned up table; Installed hex beam dumps 
    Reply  Thu Feb 13 18:00:30 2020, anchal, DailyProgress, scatter, blocked NF1811 with hex beam dump 
Message ID: 2528     Entry time: Mon Feb 10 17:41:25 2020     Reply to this: 2533
Author: anchal 
Type: DailyProgress 
Category: scatter 
Subject: Cleaned up table; Installed hex beam dumps 

Today I cleaned up the table, removed Scott's RFAM measurement setup and installed hex beam dumps on the input rejection of faraday isolators.


Recalibration

  • I measured the laser power at (50, 26) and (50, 29) when the lasers were detuned from the cavities using a power meter. Then I updated the ratio of DC readout of SN010 and SN009 with these power levels to infer reflected power in mW correctly.
  • Similarly, I measured the laser power at (9, 26) and (9, 29) right after the cavities and updated ratio with dc readout of cavity transmission photodiodes at (4, 37) and (9, 34) to infer transmitted power correctly.
  • I also adjusted horizontal tilt of periscope top mirrors at (46, 26) and (46, 29) to improve modematching to about 68% on North side and 75% on South side (numbers calculated by ratio of transmission to reflected power).

Table cleanup and new beam dumps

  • Not complete, but I removed a lot of unused optics, beam dumps, clamps etc from the table.
  • I removed the photodiodes I put in before and after PMC to measure the intensity noise in the north path and its origin (CTN:2502).
  • I dismantled Scott's RFAM measurement set up (CTN:2377) on the South path which used pick off from the input rejection of the faraday isolator at (58, 19).
  • I have now installed new hex beam dumps at the input rejection of the faraday isolators on path path at (62, 16) and (65, 27).

 

ELOG V3.1.3-